The measure is 47 less than twice that of its complement.

Answer by checkley77(12569)

2(90-x)-47+x=90
180-2x-47+x=90
-x=90+47-180
-x=-43
x=43 ans.
Proof:
2(90-43)-47+43=90
2*47-47+43=90
94-47+43=90
90=90